Emmett:

Given a polynomial equation, p(x)=x2–x–2. Find the zeros of the equation.

darkknight:

graphically or algebraically

Emmett:

Algebraically

darkknight:

Quadratic formula always works, do you know what it is?

snowflake0531:

\[ax^2+bx+c\\ ~\\a=1\\b=-1\\c=-2\] \[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

snowflake0531:

\[\frac{ -(-1) \pm \sqrt{(-1)^2-4(1)(-2)} }{ 2(1) }\]

snowflake0531:

After you simplify that you can get your two roots